Determine the concentration of hydroxyl ions in 0.4M NH4OH solution having 1)no NH4Cl 2)5.35gm of NH4Cl


ionization constant of NH4OH is 1.8*10-5

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Tejas,

1) NH4Cl---->NH4+ + Cl-

    0.4(1-x)      0.4x    0.4x

    0.4*x2/(1-x)=1.8*10-5

     x2/(1-x)=4.5*10-4

     as the value of ionization constant is less 1-x≈1

      x2=4.5*10-4

      [OH-]=2.12*10-2

2)NH4Cl completely undergoes dissociation

   5.35 g of NH4Cl is 0.1 moles

    x(0.1+x)/(0.4-x)=1.8*10-5

    x2+x(0.1+1.8*10-5)-7.2*10-4=0

    0.1+1.8*10-5≈01

    x2+0.1x-7.2*10-4=0

    x=6.745*10-3

   [OH-]=6.745*10-3